# ExPLoRA: Parameter-Efficient Extended Pre-Training to Adapt Vision Transformers under Domain Shifts

### v2

@@ 6.2 WiLDS Datasets
The WILDS Camelyon17 dataset consists of images of cancerous and non-cancerous cell tissue organized in labeled and unlabeled splits. We use the “train-unlabeled" split for pre-training ExPLoRA, and either use LoRA fine-tuning or linear probing on the training set of the labeled split. We report accuracy on the binary classification problem and compare with entries on the WILDS leaderboard which use unlabeled data. Our results in table 9 demonstrate improved performance over domain-specific methods as well as DinoV2, once again successfully bridging the domain gap.
@@
iWildcam classification requires identifying one of 182 animal species given an image. We pre-train on the training set, finding that this outperforms pre-training on the extra-unlabeled set. In table 9, we find an improvement over DinoV2 using LoRA-$r8$ PEFT. Surprisingly, the linear probing performance of the ExPLoRA suffers in comparison with DinoV2, suggesting possible loss in knowledge-transfer due to a small domain gap. Likely because natural image datasets $W_{D_{S}}$ such as ImageNet (Deng et al., 2009) used for DinoV2 already contain many images of animals.
@@
The GlobalWheat dataset consists of a wheat head object detection task, where each image of a wheat field is associated with bounding boxes on the visible wheat heads David et al. (2020; 2021). ExPLoRA extends pre-training on the training set, and then we run fine-tuning using Detectron2 code for object-detection with ViTs (Wu et al., 2019). ExPLoRA outperforms both fully pre-trained baselines from the WILDS leaderboard and strong VFMs DinoV2 and MAE on top 1 accuracy, average precision, and average recall.
6.3 Analyzing ExPLoRA
Figure 3: The mean of the principal components of the feature map outputted by each ViT block.
@@ 6.3 Analyzing ExPLoRA
Figure 4: Linear probing each patch for position (local information), across all ViT blocks.
@@
Figure 5: Linear probing each patch for classification (global information), across all ViT blocks.
@@
The key design choice of ExPLoRA is to fully train a small subset ${\mathcal{U}}\subset{\mathcal{L}}$ of the ViT, while applying low-rank updates to the remaining frozen layers ${\mathcal{L}}\setminus{\mathcal{U}}$. For parameter-efficiency, we aim to keep $|{\mathcal{U}}|\ll|{\mathcal{L}}|$ and make and informed choice of which layers to unfreeze based on their potential to improve learning during extended pre-training.
@@
We conduct an investigation on 5 models using a sample of ${\mathcal{X}}_{D_{T}}$. These models are DinoV2, D-ExPLoRA-[L]-$r64$, SatMAE, MAE, and M-ExPLoRA-[L]-$r64$. We do the following analyses: (i) PCA to measure the mean and variance of eigenvalues of patch feature vectors for each ViT block, in fig. 3 (ii) linear probing for local or global information (Darcet et al., 2023) by training logistic regression classifiers on each block’s patch feature vectors, to predict either patch position (fig. 4) or image class (fig. 5).
Findings and Unfreezing Strategy for DinoV2
@@
Our analysis reveals that the spectral properties of a block’s feature map (fig. 3) and the ability to retrieve local information from its output patch tokens (fig. 4) are correlated. The classification accuracy for position and the mean of the principal eigenvalues peak in the middle-layers of the model, suggesting that the middle blocks capture local properties of patches (e.g., texture, relative position). Meanwhile, deeper blocks focus on global semantic understanding, as shown by increased classification accuracy for image class prediction in fig. 5. Combined, these results suggest that unfreezing deeper layers, such as ${\mathcal{U}}=\{\texttt{L}\}$, allows the model to better capture global features without overfitting to local details of images of $D_{T}$. This is empirically confirmed in table 3, where linear probing accuracy correlates inversely with the mean eigenvalue of each block (i.e., block 23 > block 22 > block 0 > block 9). The attention maps in fig. 6 further support this, showing that the deeper layers focus more clearly on central objects, while earlier layers (e.g., blocks 9, 10) exhibit more diffuse attention patterns spread around the border.
Findings and Unfreezing Strategy for MAE
@@
For MAE, we see a similar, but less pronounced trend. However, MAE is only trained for reconstruction, and so retains more local information across the ViT’s layers. This is reflected by its lower patch-wise eigenvalues, higher localization accuracy, and lower global accuracies than Dino.
ExPLoRA’s Impact
@@
D-ExPLoRA preserves local information in the middle layers but also improves localization accuracy in the last few layers. Importantly, it also enhances the global information contained in the patches for deeper model layers. This indicates a better understanding of the target domain, as seen in A.2, where ExPLoRA’s attention highlights the central object more clearly.
@@ 7 Conclusion and Discussion
In this paper, we introduce ExPLoRA, a novel pre-training strategy to adapt pre-trained ViT foundation models for natural images to additional visual domains such as satellite imagery or medical data. We challenge the common paradigm of expensive pre-training from scratch for each new visual domain by offering a solution to transfer knowledge from foundation models that is both parameter-efficient and effective (even outperforming domain-specific foundation models). Our hope is that ExPLoRA enables further use of foundation models on domains other than natural images without requiring vast computational resources for pre-training.
While effective, there are many aspects of ExPLoRA that deserve further study. The strategy of fully training a small amount (or budget) of weights combines extremely well with PEFT techniques such as LoRA– we hope that future work investigates the reason behind this in further detail. Unresolved questions also include whether other parameter-efficient techniques might work better with ExPLoRA during pre-training. Further work to evaluate ExPLoRA for natural language would be valuable, as would an investigation into whether we can do away entirely with unfreezing a transformer block.
@@
As the scale of models and datasets grows exponentially, access to the computing power necessary to develop and make use of foundation models is increasingly restricted to the hands of a few organizations. This leaves many researchers in academia or smaller companies reliant on the resources of such organizations for ML research and applications. Techniques such as PEFT can alleviate this dependence and enable those with fewer computational resources to adapt, investigate, and customize models for their own needs. We hope that ExPLoRA furthers this goal, allowing ML practitioners to tailor foundation models with minimal compute, thus broadening access to powerful ML tools for critical fields like sustainability and medicine.
For example, automated analysis of satellite imagery can inform social, economic, and environmental policies, but manual curation is expensive, and pre-training models on such data has significant costs, both environmental and otherwise (see appendix C). ExPLoRA offers a more efficient way to distill knowledge from existing foundation models trained on natural images, sharply reducing costs while aiding researchers and policymakers and enabling flexible applications in downstream tasks.
@@ Acknowledgements
We include supplementary material in the following sections.
Appendix A Additional Experimental Results
@@
A.1 Results on Additional Downstream Datasets
@@
We consider a land-cover classification dataset used in Ayush et al. (2021), where each of 244,471 training and 55,529 validation images are paired with one of 66 land cover classes obtained by the USDA’s National Agricultural Imagery Program. In table 10, we first demonstrate similar performance between both natural-image backbones (rows 4 and 5), which surprisingly outperform SatMAE, which is pre-trained on fMoW-RGB. We use ExPLoRA to pre-train from DinoV2 to the training set of this dataset (without labels). Our results (row 6) demonstrate comparable performance, suggesting that for this dataset, domain-specific knowledge may not be highly relevant to successfully solve the task.
@@
The dataset contains 27,000 13-band satellite images of 10 classes Helber et al. (2019), sourced from Sentinel-2. For ExPLoRA, we don’t pre-train on the training set of this dataset, and instead use LoRA fine-tuning starting with the pre-trained weights learned in row 8 of table 4. We demonstrate improved performance over DinoV2, and match the performance achieved by the domain-specific SatMAE which was fully pre-trained on fMoW-Sentinel, and fully fine-tuned on EuroSAT (table 11). This demonstrates the successful use of our extended pre-trained model on further downstream datasets.
A.2 Attention Map Visualizations
To aid our analysis in section 6.3, we visualize attention scores for different ViT blocks across multiple models, including DinoV2, D-[L]-$r64$ (i.e. the last row of table 3), the second and third rows of table 3, MAE, SatMAE, and M-[L]-$r64$. These visualizations are shown in fig. 6 for 3 different images from the validation set of fMoW-RGB. Since our models are trained without registers, we truncate attention scores more than 5 standard deviations away from the mean, thus removing artifact attention scores with unusually high values on background patches (Darcet et al., 2023).
The visualizations in fig. 6 further support the analysis in section 6.3. For the Dino models, the attention scores of block 9-10 are diffuse and spread around the central object of the image, with quite a few border pixels highlighted. Conversely, the attention scores of the final layers are concentrated more towards the central object. These visualizations further suggest that the middle layers focus on capturing local properties of the images such as texture, while the final layers capture global semantic information such as object-ness. Interestingly, the initial blocks for the Dino models display sparse attention patterns with spikes on seemingly random patches. This might suggest a form of caching to aid the computation of deeper layers that will extract local or global information.
For the MAE models, we see that the original MAE (pre-trained on natural images) seem to highlight more border pixels in the final layers of the ViT. Post extended pre-training with ExPLoRA, the final layers concentrate attention scores on the central object, more closely resembling the patterns of SatMAE (which was fully pre-trained on satellite images). ExPLoRA is thus able to successfully transfer knowledge from its initialized source-domain weights $W_{D_{S}}$ to serve as a foundation model $W_{D_{T}}^{\ast}$ on the new target domain $D_{T}$.
A.3 Convergence and Data Efficiency
@@
One hypothesis for the effectiveness of pairing unfreezing blocks with LoRA tuning is that the low-rank updates to the ViT backbone “nudge" the sequence of embedded visual tokens from $D_{S}$ to those representing $D_{T}$, which then enables the unfrozen ViT block to efficiently compress global information from the new domain.
A.4 Impact of MAE Decoder Rank
@@
As outlined in section 5, we initialize the MAE decoder ${\mathcal{L}}_{D}$ with pre-trained weights $W_{D_{S}}$ from He et al. (2022), keeping all decoder weights (except layer norm) frozen during extended pre-training on ${\mathcal{X}}_{D_{T}}$. We apply LoRA with rank $r^{\prime}$ to the $Q,V$ weights of the attention layers in the decoder ${\mathcal{L}}_{D}$, while unfreezing 1-2 blocks ${\mathcal{U}}$ in the ViT encoder ${\mathcal{L}}$ and applying LoRA with rank $r$ to the remaining layers ${\mathcal{L}}\setminus{\mathcal{U}}$ (algorithm 1).
We evaluate ExPLoRA with M-$\tt{[1,L]}$-$r64$ on fMoW-Sentinel, using a fixed encoder LoRA rank $r=64$, unfreezing blocks ${\mathcal{U}}=\{\tt{1,L}\}$, and varying the decoder rank $r^{\prime}$ We then fine-tune the resulting model with LoRA $r=8$ and measure the highest top 1 accuracy on the validation set of fMoW-Sentinel. Results in table 12 show that increasing $r^{\prime}$ up to 32 improves fine-tuning performance, but performance declines when $r^{\prime}=64$, dropping $\downarrow 0.94\%$. This suggests that balancing the unfrozen parameters between the ViT encoder ${\mathcal{L}}$ (used for fine-tuning) and the MAE decoder ${\mathcal{L}}_{D}$ (discarded post pre-training) is crucial. Larger $r^{\prime}$ may improve the decoder’s ability without benefiting the learned representations of ${\mathcal{L}}$. This issue doesn’t arise in DinoV2, as the Dino-iBOT shared head is fully trained since it isn’t provided by Oquab et al. (2023).
A.5 Impact of ViT backbone size

### v3

@@ 6.2 WiLDS Datasets
The WILDS Camelyon17 dataset consists of images of cancerous and non-cancerous cell tissue. We use the “train-unlabeled" split for pre-training ExPLoRA, and either use LoRA fine-tuning or linear probing on the training set of the labeled split. We report accuracy on the binary classification problem and compare with entries on the WILDS leaderboard which use unlabeled data. Our results in table 9 demonstrate improved performance over domain-specific methods as well as DinoV2, once again successfully bridging the domain gap.
@@
iWildcam classification requires identifying one of 182 animal species given an image. We pre-train on the training set, finding that this outperforms pre-training on the extra-unlabeled set. In table 9, we find an improvement over DinoV2 using LoRA-$r8$ PEFT. Surprisingly, the linear probing performance of the ExPLoRA suffers in comparison with DinoV2, suggesting possible loss in knowledge-transfer due to a small domain gap. Likely because natural image datasets $W_{S}$ such as ImageNet (Deng et al., 2009) used for DinoV2 already contain many images of animals.
@@
The GlobalWheat dataset consists of an object detection task, where images of wheat fields are associated with bounding boxes on the visible wheat heads (David et al., 2020, 2021). ExPLoRA extends pre-training on the training set, and then we fine-tune using Detectron2 code for object-detection with ViTs (Wu et al., 2019). ExPLoRA outperforms both fully pre-trained baselines from the WILDS leaderboard and strong VFMs DinoV2 and MAE on top 1 accuracy, average precision, and average recall.
7 Further Analysis
Figure 3: The mean of the eigenvalues of the feature map outputted by each ViT block.
Figure 4: The variance of the eigenvalues of the feature map outputted by each ViT block.
@@ 7 Further Analysis
Figure 5: Linear probing patches for position (local information), across all ViT blocks.
@@
Figure 6: Linear probing patches for classification (global information), across all ViT blocks.
@@
The key design choice of ExPLoRA is to constrain the trainable parameter space during pre-training via $\Theta({\mathcal{U}},r)$ in eq. 5, where ${\mathcal{U}}\subset{\mathcal{L}}$ layers of the ViT undergo full-rank training while the remaining frozen layers ${\mathcal{L}}\setminus{\mathcal{U}}$ receive low-rank updates. For parameter-efficiency, we wish to keep $|{\mathcal{U}}|\ll|{\mathcal{L}}|$ and make an informed choice of which layers to unfreeze based on their potential to improve representation learning during extended pre-training.
@@
We conduct an investigation on 5 models using a sample of ${\mathcal{X}}_{D_{T}}$. These models are DinoV2, D-ExPLoRA-[L]-$r64$, SatMAE, MAE, and M-ExPLoRA-[L]-$r64$. We do the following analyses: (i) PCA to measure the mean and variance of eigenvalues of patch feature vectors for each ViT block, in figs. 3 and 4 (ii) linear probing for local or global information (Darcet et al., 2023) by training logistic regression classifiers on each block’s patch feature vectors, to predict either patch position (fig. 5) or image class (fig. 6).
@@
Findings and Unfreezing Strategy for DinoV2: Our analysis reveals that the spectral properties of a block’s feature map (fig. 3) and the ability to retrieve local information from its output patch tokens (fig. 5) are correlated. The classification accuracy for position and the mean and variance of the principal eigenvalues peak in the middle-layers of the model, suggesting that the middle blocks capture fine-grained local properties of patches (e.g., texture, relative position). Meanwhile, deeper blocks focus on global semantic understanding, as shown by increased classification accuracy for image class prediction in fig. 6 and lower variance in feature map eigenvalues in fig. 4. Combined, these results suggest that unfreezing deeper layers, such as ${\mathcal{U}}=\{\texttt{L}\}$, allows the model to better capture global features without overfitting to local details of images of $T$. This is empirically confirmed in table 3, where linear probing accuracy correlates inversely with the mean eigenvalue of each block (i.e., block 23 > block 22 > block 0 > block 9). The attention maps in fig. 9 further support this, showing that deeper layers focus more clearly on central objects, while earlier layers (e.g., blocks 9, 10) exhibit diffuse attention patterns spread around the border.
@@
Findings and Unfreezing Strategy for MAE: For MAE, we see a similar, but less pronounced trend. However, MAE is only trained for reconstruction, and so retains more local information across the ViT’s layers. This is reflected by its lower patch-wise eigenvalues, higher localization accuracy, and lower global accuracies than Dino.
@@
ExPLoRA’s Impact: D-ExPLoRA preserves local information in the middle layers but also improves localization accuracy in the last few layers. Importantly, it also enhances the global information contained in the patches for deeper model layers. This indicates a better understanding of the target domain, as seen in B.8, where ExPLoRA’s attention highlights the central object more clearly.
@@ 8 Conclusion and Discussion
In this paper, we introduce ExPLoRA, a novel pre-training strategy to adapt pre-trained ViT foundation models for natural images to additional visual domains such as satellite imagery or medical data. We challenge the common paradigm of expensive pre-training from scratch for each new visual domain by offering a solution to transfer knowledge from foundation models that is both parameter-efficient and effective (even outperforming domain-specific foundation models). Our hope is that ExPLoRA enables further use of VFMs on domains other than natural images without requiring vast computational resources for pre-training.
While effective, many aspects of ExPLoRA deserve further study. The strategy of fully training a small budget of weights combines well with PEFT techniques like LoRA– understanding this further would be valuable. Future work might explore whether other parameter-efficient techniques could improve ExPLoRA during pre-training more effectively than unfreezing blocks. Lastly, an investigation of ExPLoRA on large language models would be valuable.
@@
As the scale of models and datasets grows exponentially, access to the computing power necessary to develop and use foundation models is increasingly restricted to the hands of a few organizations. This leaves many researchers in academia or smaller companies reliant on the resources of such organizations for ML research and applications. Techniques such as PEFT can alleviate this dependence and enable those with fewer computational resources to adapt, investigate, and customize models for their own needs. We hope that ExPLoRA furthers this goal, allowing ML practitioners to tailor foundation models with minimal compute, thus broadening access to powerful ML tools for critical fields like sustainability and medicine.
For example, automated analysis of satellite imagery can inform social, economic, and environmental policies, but manual curation is expensive, and pre-training models on such data has significant costs, both environmental and otherwise (appendix D). ExPLoRA offers a more efficient way to create new foundation models for different visual domains via distilling knowledge from existing foundation models trained on natural images. This can sharply reduce costs, aid researchers and policymakers, and enable flexible downstream applications.
@@ Acknowledgements
We include supplementary material in the following sections. We provide additional contextualization and comparison with recent related work in appendix A, additional experimental results and ablations in appendix B, training and hyperparameter details in appendix C, and an overview of the environmental impact of our work in appendix D.
Appendix A Further Contextualization with Related Work
Here, we expand upon ExPLoRA’s differences with recent related work.
A.1 Comparison with Geospatial Domain Adaptation
Recent work has explored both continual pre-training (GFM) (Mendieta et al., 2023) and parameter-efficient domain adaptation (GDA) (Scheibenreif et al., 2024) for satellite imagery. We compare these approaches with ExPLoRA in table 10.
GFM GDA ExPLoRA
[44] [56] (Ours)
Parameter-efficient
Training objective MAE MAE Any
Arch. preservation
Fine-tuning Any PEFT Only LoRA Any PEFT
Table 10: Differences between prior geospatial domain adaptation methods and ExPLoRA
ExPLoRA differs from these approaches in several key aspects. Unlike GFM which trains the full backbone, ExPLoRA achieves superior performance with only a fraction of trainable parameters. While GDA is also parameter-efficient, it requires non-mergeable scaling vectors that induce inference latency and modify the ViT, whereas ExPLoRA’s LoRA adapters can be merged into the ViT’s weights. Additionally, ExPLoRA extends beyond MAE architectures (supporting DinoV2 and others) and allows flexible configurations between pre-training and fine-tuning, including varying LoRA ranks or using different PEFT methods, which GDA doesn’t support out-of-the-box.
We also demonstrate ExPLoRA’s broader applicability through experiments on larger datasets (fMoW-RGB, fMoW-Sentinel, which have 400k-800k images vs 90k images in FireRisk (Shen et al., 2023), the largest dataset used in GDA) and domains beyond remote sensing (i.e. WiLDS). Our analysis in section 7 provides insights into block-wise information encoding, offering practitioners a systematic approach for block selection during PEFT– a unique feature not present in prior work.
A.2 Comparison with Unsupervised Domain Adaptation
UDA ExPLoRA
Source data Labeled None
Source knowledge Data Weights
Target data Unlabeled Unlabeled
Label constraints $Y_{T}\subseteq Y_{S}$ None
Table 11: Differences between UDA and ExPLoRA
Unsupervised domain adaptation (UDA) enables models to generalize to unseen domains (Kang et al., 2019; Oren et al., 2019; Singhal et al., 2023; Khanna et al., 2023). Traditional UDA assumes:
  1. (i)
    ${\mathcal{Y}}_{S}=\{{\mathbf{x}}_{i},{\mathbf{y}}_{i}\}_{i=1}^{N^{\prime}}\sim p_{S}({\mathbf{x}},{\mathbf{y}})$, a labeled source domain dataset
  2. (ii)
    ${\mathcal{X}}_{T}=\{{\mathbf{x}}_{i}\}_{i=1}^{N}\sim p_{T}({\mathbf{x}})$, an unlabeled target domain dataset
  3. (iii)
    $Y_{T}\subseteq Y_{S}$, constraining the label-set of $T$ with respect to $S$
Common UDA benchmarks like Office-Home (Venkateswara et al., 2017) and VisDA-2017 (Peng et al., 2017) follow this setup (Xu et al., 2021; Sun et al., 2022; Yang et al., 2023; Zhu et al., 2023).
ExPLoRA’s setting in section 4 is different: we only require weights $W_{S}$ from unsupervised pre-training on $p_{S}({\mathbf{x}})$, without source data access or label set restrictions. This enables adaptation across wider domain shifts (e.g., ImageNet to multi-spectral satellite imagery, section 6.1.3). Thus, rather than competing with UDA methods, ExPLoRA can complement them by providing a better initialization than standard natural-image pre-training (as seen empirically in section B.7).
Appendix B Additional Experimental Results
@@
B.1 Results on Additional Downstream Datasets
@@
We consider a land-cover classification dataset used in (Ayush et al., 2021), where each of 244,471 training and 55,529 validation images are paired with one of 66 land cover classes obtained by the USDA’s National Agricultural Imagery Program. In table 12, we first demonstrate similar performance between both natural-image backbones (rows 4 and 5), which surprisingly outperform SatMAE, which is pre-trained on fMoW-RGB. We use ExPLoRA to pre-train from DinoV2 to the training set of this dataset (without labels). Our results (row 6) demonstrate comparable performance, suggesting that for this dataset, domain-specific knowledge may not be highly relevant to successfully solve the task.
@@
The dataset contains 27,000 13-band satellite images of 10 classes (Helber et al., 2019), sourced from Sentinel-2. For ExPLoRA, we don’t pre-train on the training set of this dataset, and instead use LoRA fine-tuning starting with the pre-trained weights learned in row 8 of table 6. We demonstrate improved performance over DinoV2, and match the performance achieved by the domain-specific SatMAE which was fully pre-trained on fMoW-Sentinel, and fully fine-tuned on EuroSAT (table 13). This demonstrates the successful use of our extended pre-trained model on further downstream datasets.
B.2 The Importance of Extended Pre-training
Figure 7: Fine-tuning accuracy versus total compute (measured in GPU-hours). Total compute includes both pre-training (if applicable) and fine-tuning phases. Along with the label for each method in the legend, we include (#pre-training params, #fine-tuning params).
To evaluate ExPLoRA’s effectiveness, we analyze how its performance scales with computational resources. Specifically, we investigate two key questions: First, given a fixed compute budget, what is the optimal allocation between extended pre-training and fine-tuning? Second, for a fixed parameter budget, does investing compute in extended pre-training provide advantages over standard fine-tuning approaches?
We address these questions in fig. 7, focusing on DinoV2 models running on NVIDIA-A4000 GPUs. We evaluate D-ExPLoRA-$\tt{[L]}$-$r64$ for different lengths of pre-training (50k, 100k, 150k, and 200k iterations), corresponding to 24, 48, 72, and 96 GPU-hours of extended pre-training respectively. Each checkpoint undergoes LoRA-$r8$ fine-tuning. We compare against three baselines: (i) Direct LoRA-$r8$ fine-tuning on DinoV2 weights (ii) Fine-tuning DinoV2 with block 24 unfrozen and LoRA-$r64$ (matching ExPLoRA’s parameter budget) (iii) Fine-tuning DinoV2 with blocks 0, 1, 23, 24 unfrozen and LoRA-$r64$ (55.8M parameters vs ExPLoRA’s 18.7M).
Results in fig. 7 demonstrate that ExPLoRA’s extended pre-training achieves a $\uparrow 1.0\%$ improvement in maximum top-1 fine-tuning accuracy within the same total compute budget (320 GPU hours). Notably, even increasing the parameter budget during fine-tuning fails to match this performance. While additional pre-training iterations beyond 50k improve initial fine-tuning accuracy, they have minimal impact on the final accuracy ceiling, highlighting ExPLoRA’s computational efficiency.
B.3 Convergence and Data Efficiency
@@
One hypothesis for the effectiveness of pairing unfreezing blocks with LoRA is that low-rank updates to the ViT backbone “nudge" the sequence of embedded visual tokens from $S$ to those representing $T$, which then enables the unfrozen ViT block to efficiently compress global information from the new domain.
B.4 Impact of MAE Decoder Rank
@@
As outlined in section 5, we initialize the MAE decoder $g_{\psi}$ (with its ${\mathcal{L}}_{D}$ transformer layers) with pre-trained weights $W_{S}$ from He et al. (2022), keeping all decoder weights (except layer norm) frozen during extended pre-training on ${\mathcal{X}}_{T}$. We apply LoRA with rank $r^{\prime}$ to the $Q,V$ weights of the attention layers in the decoder ${\mathcal{L}}_{D}$, while unfreezing 1-2 blocks ${\mathcal{U}}$ in the ViT encoder ${\mathcal{L}}$ and applying LoRA with rank $r$ to the remaining layers ${\mathcal{L}}\setminus{\mathcal{U}}$ (algorithm 1).
We evaluate ExPLoRA with M-$\tt{[1,L]}$-$r64$ on fMoW-Sentinel, using a fixed encoder LoRA rank $r=64$, unfreezing blocks ${\mathcal{U}}=\{\tt{1,L}\}$, and varying the decoder rank $r^{\prime}$. We then fine-tune the resulting model with LoRA $r=8$ and measure the highest top 1 accuracy on the validation set of fMoW-Sentinel. Table 14 shows that increasing $r^{\prime}$ up to 32 improves fine-tuning performance, which then declines by $\downarrow 0.94\%$ for $r^{\prime}=64$. This suggests that balancing the unfrozen parameters between the ViT encoder $f_{\theta}$ (used for fine-tuning) and the MAE decoder $g_{\psi}$ (discarded post pre-training) is crucial. Larger $r^{\prime}$ may improve the decoder’s ability without benefiting the learned representations of $f$. This issue doesn’t arise in DinoV2, as the Dino-iBOT shared head is fully trained since it isn’t provided by Oquab et al. (2023).
B.5 Impact of ViT backbone size